OpenStudy (anonymous):

Which fraction is not equivalent to -(5/8)? A. (-5/8) B. (5/-8) C. (-5/-8) D. all of these

OpenStudy (anonymous):

\[\frac{-5}{8} = \frac{-1}{1} \cdot \frac{5}{8} = -\left(\frac{5}{8}\right)\] \[\frac{5}{-8} = \frac{1}{-1} \cdot \frac{5}{8} = -\left(\frac{5}{8}\right)\] \[\frac{-5}{-8} = \frac{-1}{-1} \cdot \frac{5}{8} = \left(\frac{5}{8}\right) \text{ Because the negative 1s cancel}\] So which do you think it is?

OpenStudy (anonymous):

-(5/8) is implying that it is a negative fraction. Let's go over the choices: Choice A means that (-5/8) which means it's a negative fraction. So it is equal with the problem. Choice B means (5/-8). Now when there is one negative with the fraction, then it turns the whole fraction into negative. Choice C is (-5/-8) which means there are 2 negatives. Now when two negatives are together, they cancel out and it becomes a positive. Choice D is implying that it is all of them, which is not true. Now based on my explanation, what do you think it is? Remember it's asking you to choose the one that is not the same as the problem.